OpenStudy (anonymous):

miss moneypenny inherits £330 from a secret agent she decides to save some of the money and spend the rest the ratio of savongs to spending money is 7:4 how much does she save and how much does she spend savings: spendong money:

OpenStudy (anonymous):

330/11 = 30. 30*7/30*4 =210/120

OpenStudy (anonymous):

Savings 210 spending 120

OpenStudy (anonymous):

thnx cn i ask u more

OpenStudy (anonymous):

Yeah you can ask as many questions as you like, if I can't or don't solve them someone else will.

OpenStudy (anonymous):

arthur is organising his collection of DVDS .He has 135 altogether. the ratio of comedy films to aztion to horror is 2:8:5 howmany of each DVD does he have? comedy action horror

OpenStudy (anonymous):

So add up the ratios 2 + 8 + 5 = 15 135 / 15 = 9 = 2(9):8(9):5(9) =18:72:45

OpenStudy (anonymous):

whats comdey whats action and whats horror

OpenStudy (anonymous):

???

OpenStudy (anonymous):

I have gave you the answer 18:72:45 If they are listen in order of ratio then it would be comedy 18, action 72, horror 45.

OpenStudy (anonymous):

cn i ask u more plz :) xxx

OpenStudy (anonymous):

Only if you attempt to solve it first yourself.